# Relative Distance Guided Dynamic Partition Learning for Scale-Invariant UAV-View Geo-Localization

### v2

@@ IV-A Datasets and Evaluation Metrics
Datasets. We perform extensive experiments based on the University-1652 [10] dataset for the following reasons: 1) the University-1652 contains diverse scenarios, including 1652 university buildings worldwide, where each building comprises 1 satellite-view image and 54 drone-view images. The training set includes 701 buildings of 33 universities, and the other 951 buildings of the remaining 39 universities are classified as the test set. 2) drone-view images exhibits considerable scale variation, with distances raning from 123.5m to 256m (i.e., $H_{D}\in[123.5m,256m]$), supporting the verification of the anti-scale robustness of our DGDPL. Referring to LPN [24], the test set is equally divided into three parts according to the drone-view distance: Long, Middle, and Short. Then, comparing satellite-view images with UAV-view images, we set the value of $H_{S}$ to 189.75m, which corresponds to the midpoint of the UAV distance range.
Figure 5: The data augmentation strategy to extend the distance range of drone-views from the visual perspective. (a) Making the distance seem farther: given a drone image, mirror a strip $\Delta$P pixels wide around the input image and resize it to initial resolution; (b) Making the distance seem closer: given a drone image, crop out a square ring $\Delta$P pixels wide around the input image and resize it to initial resolution.
To further expand the distance range of drone-views, we design a data augmentation strategy to simulate more extreme image scales. As shown in Fig. 5, we introduce a parameter $\Delta P$ to adjust image scales (equivalent to UAV distance), which can be expressed as:
$x_{D}^{\Delta P}=\mathcal{F}_{aug}(x_{D};\Delta P)$ (16)
where $\mathcal{F}_{aug}$ denotes the process of data augmentation, $x_{D}$ and $x_{D}^{\Delta P}$ denote the input image and synthesized image, respectively. ($\Delta P$$>$0) means that making the distance of drone-view seem farther (see Fig. 5(a)), while ($\Delta P$$<$0) means that making the distance of drone-view seem closer (see Fig. 5(b)). $|\Delta P|$ denotes the degree of scale adjustment. Then, we roughly calculate the distance of synthesized image $x_{D}^{\Delta P}$:
$H_{D}(x_{D}^{\Delta P})=H_{D}(x_{D})+\lambda_{aug}*\Delta P$ (17)
where $\lambda_{aug}$=0.7 is the empirical value, estimated by the visual scale and distance of the existing drone-view images.
By applying the aforementioned strategy to Long and Short sets, we generate test sets with extreme scales, named as Long($\Delta P$$>$0) and Short($\Delta P$$<$0), respectively. Note that drone-views within the Long($\Delta P$$>$0) exhibit greater distances, whereas those within the Short($\Delta P$$<$0) have shorter distances.
@@
ResNet [80] CVPR’16 ResNet-50 60.16 64.98 37.62 43.36 23.12 28.88 40.30 45.74
@@
DGDPL ViT-S 88.08 89.77 75.12 78.33 78.15 81.34 80.45 83.14
IV-B Model Settings and Training Details
@@
Model Settings. For the default DGDPL framwork, we set the feature map upsampling factor $\lambda$=4, the number of partitions to $N_{SPS}$=4, and partition scaling factor $\alpha$=3.3. For the SRGS, we adopt Chebyshev distance to generate coordinate map $CM$, and set the heatmap segmentation threshold $\delta$=0.5. For the University-1652 dataset, we roughly set drone-views distance $H_{D}\in[123.5m,256m]$ and satellite-views distance $H_{S}$=$189.75m$, respectively.
@@
Training Details. By default, we resize each input image to 512$\times$512 for both training and testing phases, and employ the ResNet-50 [80] to extract visual features. In training, we set the initial learning rate of 0.001 for the backbone, and the rest of learnable parameters are set to 0.0001. Stochastic gradient descent (SGD) are adopted with batch size 4, momentum 0.9, and weight decay 0.0005 for optimization. The model are trained for 120 epochs, and the learning rate is decayed by 0.1 after 80 epochs. We also employ random horizontal image flipping as data augmentation. During testing phase, Euclidean distance are adopted to measure the similarity between the query images and candidate images in the gallery set. Experiments are performed on an NVIDIA RTX 3090 GPU with 24 GB of memory.
@@
Figure 6: Image retrieval results obtained with DGDPL, LPN and ResNet. Specifically, we present the correspondingly numbered image retrieval outcomes from the Test set (a), Long($\Delta P$=100)(b), and Short($\Delta P$=-100)(c), respectively.
IV-C Comparison with State-of-the-art
@@
As shown in Table I, we compare the proposed DGDPL with other methods on University-1652. Except for the standard test set provided, we employ Eq. (16) to generate two supplementary test sets, named Long($\Delta P$=100) and Short($\Delta P$=-100), which exhibit more extreme distance of drone-views. In other words, Long($\Delta P$=100) and Short($\Delta P$=-100) have more significant cross-view scale differences. The quantitative results are divided into two groups, ResNet-based and ViT-based methods. On the ResNet-based track, DGDPL achieves 84.19% R@1 and 86.36% AP on default test set, 65.39% R@1 and 69.41% AP on Short($\Delta P$=-100) set and 63.42% R@1, and 68.05% AP on Long($\Delta P$=100) set, respectively. On the default test set, DGDPL achieves competitive performance with recent SPDL [55] and clearly outperforms other solutions, such as $D^{2}$-GeM [14] and Sample4Geo [38]. Moreover, it is remarkable that our DGDPL yields superior results on both the Short($\Delta P$=-100) and Long($\Delta P$=100) compared with partition-based LPN [19] and SDPL [55]. Specifically, compared with LPN, DGDPL improves R@1 from 47.24%, 35.57% to 65.39% (+18.15%), 63.42% (+27.85%), respectively. On the ViT-based track, DGDPL is compatible with the ViT framework and has surpassed the sate-of-the-art methods on Short($\Delta P$=-100) and Long($\Delta P$=100), achieving an average R@1 and AP of 80.45% and 83.14%. Compared with ViT [79] and FSRA, DGDPL significantly improves retrieval accuracy in all scenarios, with an average R@1 improvement of 10.59% and 6.37%. This shows that DGDPL is robust to scenes with inconsistent cross-view scales.
@@
Moreover, we visualize some retrieved results in Fig. 6. To prove the anti-scaling robustness of DGDPL, we provide same numbered results generated by baseline [80] and LPN [24] on different test sets. For the original test set, both LPN and DGDPL are able to retrieve the correct satellite labels. As shown in Fig. 6(b) and (c), when the drone-view images undergo more drastic scale variations, it can be observed that our DGDPL still achieves correct results. This benefits from the fact that DGDPL explicitly aligns contents of partition pairs, thus mining robust contextual information. For example, as shown in Group 3 of Fig. 6(b), while the target in the satellite-view is discernible, the target size in the drone-view is notably smaller compared to that in the satellite-view. This discrepancy leads to a misalignment of features extracted by LPN with fixed templates, resulting in inaccurate retrieval results. On the other hand, as shown in Group 2 of Fig. 6(c), drone-view with shorter distance implies less environmental information and larger scale differences of cross-views. Compared to LPN, DGDPL can dynamically adjust partition sizes to boost the alignment of local features, but also refine part-level features, yielding accurate retrieval results.
IV-D Ablation Studies
To verify the effectiveness of DGDPL, we conduct extensive ablation studies, including:
  • To validate the robustness of DGDPL in scale-inconsistent scenarios and its adaptability to diverse backbone networks;
  • To prove the effectiveness of the proposed square partition strategy, distance-guided adjustment strategy and saliency-guided refinement strategy;
  • To Explore the effect of hyper-parameters on model performance, including number of partitions $N_{SPS}$, partition scaling factor $\alpha$ and input image size.
Robustness of DGDPL to scale-inconsistent scenarios. As a hard-partition-based method, our DGDPL can alleviate the dependence on scale consistency of cross-views by dynamically adjusting the size of partition pairs. To fully verify the robustness of DGDPL for scale-inconsistent scenarios, we first synthesize a series of test sets with closer and longer distance of drone-views, named Short($\Delta P$$<$0) and Long($\Delta P$$>$0), where a larger $|\Delta P|$ represents a larger UAV-view scale variation. Then we compare the proposed DGDPL with representative partition-based methods, such as LPN [24] and FSRA [19]. The results are shown in Table II, divided into ResNet50-based track and ViT-based track. The performance degradation of each method on different datasets is shown in Fig. 7. On the ResNet50-based track, DGDPL shows the best retrieval accuracy, with significantly less performance degradation than other partition-based methods. In particular, compared with LPN and SDPL, our DGDPL improves average R@1 from 56.53%, 65.02% to 70.40%(+13.87%, +5.38%), respectively. Furthermore, the integration of DGAS into the DGDPL framework led to a notable improvement in retrieval accuracy in extreme scenarios, with a 5.16% improvement in the average R@1. This highlights that DGAS facilitates alignment of fine-grained features from multiple views. On the ViT-based track, DGDPL also achieves optimal retrieval accuracy across datasets of all scales, with an average R@1 of 79.46%. Compared with FSRA and LPN, DGDPL improves average R@1 by 5.66% and 4.37%. This fully demonstrates the robustness of DGDPL to cross-view scale variations, as well as the compatibility with various backbone networks. Moreover, Fig. 7 proves that DGAS can alleviate the performance degradation of DGDPL caused by cross-view scale inconsistency without additional learnable parameters. It is worth noting that the performance degradation of ViT-based methods is generally lower than that of ResNet-based methods. We conclude that the ViT framework is capable of capturing long-range feature dependencies, thereby contributing to its robustness against scale changes. As depicted in Fig. 77(c), in contrast to ResNet-based methods, the performance degradation of ViT methods is comparable in Short($\Delta P$$<$0), whereas the baseline exhibits the minimal degradation. This phenomenon can be primarily attributed to the robust feature extraction capability of the ViT, coupled with the scarcity of effective features in Short($\Delta P$$<$0) scenarios.
@@ IV-D Ablation Studies
TABLE II: ablation study on robustness of DGDPL to scale-inconsistent scenarios. we reported the R@1 accuracy for baseline, LPN, FSRA, SDPL and DGDPL on various test sets. The experimental results are divided into two groups: Resnet50-based at the top and ViT-based at the bottom. best and second best performance are in red and blue colors, respectively.
@@
Figure 7: Performance degradation of different methods on Short($\Delta P\leq 0$) and Long($\Delta P\geq 0$). Vertical axis: magnitude of decrease in R@1. Horizontal axis: degree of distance adjustment $|\Delta P|$. (a) and (b) show the results of ResNet-based methods, while (c) and (d) show the results of ViT-based methods.
@@
Effectiveness of each component. The core components of DGDPL are SPS, DGAS and SGRS. We construct five derived models to validate the effectiveness of each strategy, and results are shown in Table III. Firstly, fine-grained partitions generated by SPS contain both local and global features, thus learning rich contextual information. Compared to the baseline, SPS provides +19.76% R@1, +21.60% R@1 and +24.88% R@1 on three test sets. When DGAS is combined with SPS, the model’s accuracy is notably enhanced in scenarios involving scale inconsistency, while maintaining stable performance in the standard test set. For instance, upon removing DGAS from DGDPL, the R@1 of model(SPS+SGRS) decreases by 0.65%, 5.31%, and 12.29% on Standard test set, Short($\Delta P$=-100) and Long($\Delta P$=100), respectively. We emphasize that drone-views with short distances contain incomplete depictions of targets and limited environmental information. consequently, it is plausible that the performance gain of the DGAS is lower in the Short set compared to the Long set. To further improve the retrieval accuracy of DGDPL, we introduce SGRS to refine salient regions of partitions. It can be observed that SGRS boosts the retrieval accuracy of DGDPL in various scenarios, yielding 3.50%, 5.64% and 2.71% R@1 improvement over the model(SPS+DGAS).
@@
TABLE III: performance comparison of DGDPL framework with different components. $\checkmark$ indicates that the module is applied.
Model strcuture Standard Test set Short($\Delta P$=-100) Long($\Delta P$=100)
SPS DGAS SGRS R@1 AP R@1 AP R@1 AP
@@
$\checkmark$ $\checkmark$ $\checkmark$ 84.19 86.36 65.39 69.41 63.42 68.05
Variants of SGRS. As mentioned above, SGRS segments part-level features based on heatmaps to obtain global features, salient features and background features, which are matched independently. To fully verify the effectiveness of SGRS, we construct a variety of SGRS variants, including different output combinations as well as coordinate map $CM$, as shown in Fig. 8. Table IV shows the experimental results. Regardless of the specific combination of output modes employed by the SGRS to refine partitions, it consistently outperforms the model without SGRS (i.e., only $g_{i}$). On the other hand, compared to Euclidean distance and Manhattan distance, the utilization of Chebyshev distance within SGRS yields an optimal gain, surpassing that of Manhattan distance by a notable margin. This phenomenon can be attributed to the fact that the $CM$ generated by the Chebyshev paradigm exhibits a structural similarity to the square partition strategy employed. Therefore, it is within the realm of expectation that the model based on Euclidean distance yields comparable results. Since the feature segmentation process of SGRS is unsupervised, the injected $CM$ can assist SGRS in judging the importance of features and classifying neighborhood features into the same class. Upon removal of the $CM$ from SGRS, the average R@1 and AP metrics of the DGDPL decreased by 6.37% and 5.72%.
@@
TABLE IV: ablation study on different variants of SGRS. The experiment aims to explore the impact of different feature combinations and coordinate maps on performance.
@@
partition scaling factor $\alpha$. Referring to Eq (4), our DGAS can adjust the degree of variation in partition size by modifying $\alpha$. An improperly choice for $\alpha$ may result in a misalignment of content between cross-view partition pairs. We set a series of $\alpha$ for experiments, and the results are shown in Table V. It can be observed that when $\alpha$ is small, the size of partition pairs of DGDPL is almost of the same size, yielding favorable outcomes in the Middle set but suffering from significant performance degradation in extreme scenarios. When $\alpha$ increases from 0.0 to 3.3, the retrieval accuracy of DGDPL continues to rise in the Short($\Delta P\leq 0$) and Long($\Delta P\geq 0$) sets. In particular, the R@1 of our method improves from 17.84% and 34.93% to 24.84%(+7.00%) and 55.64%(+20.71%) on Short($\Delta P$=-150) set and Long($\Delta P$=+150) set, respectively. The average R@1 of DGDPL also increases from 63.83% to 70.40%(+6.57%). This verifies that DGAS can enhance the robustness of our model against scale changes. As $\alpha$ continues to increase, the retrieval accuracy of DGDPL exhibits a slight decline across various Short($\Delta P\leq 0$) scenes. Considering the average performance of the model, we set $\alpha$=3.3 as the default setting.
TABLE V: performance comparison of DGDPL with various partition scaling factor $\alpha$. we reported the R@1 accuracy on three types of datasets, including Short($\Delta P\leq 0$) set, Middle set and Long($\Delta P\geq 0$) set.

### v3

@@ IV-A Datasets and Evaluation Metrics
Datasets. The main experiments are performed based on the University-1652 [8] dataset for the following reasons: 1) University-1652 contains diverse scenarios, including 1652 university buildings worldwide, where each building comprises 1 satellite-view image and 54 drone-view images. The training set includes 701 buildings of 33 universities, and the other 951 buildings of the remaining 39 universities are classified as the test set. 2) drone-view images exhibits considerable scale variation, with height ranging from 123.5m to 256m (i.e., $H_{D}\in[123.5,256]$), supporting the verification of the anti-scale robustness of SaLPN. Referring to LPN [25], the test set is equally divided into three parts according to the shooting height of drone-view: Long, Middle, and Short. Then, comparing satellite-view images with drone-view images, we set the value of $H_{S}$ to 189.75m, which corresponds to the midpoint of the drone height range.
@@
Instance loss [8] ACMMM’20 ResNet-50 60.16 64.98 37.62 43.36 23.12 28.88 40.30 45.74
@@
SaLPN ViT-S 88.19 89.79 73.45 76.91 77.44 80.69 79.69 82.46
To further expand the height range of drone-views, we design a data augmentation strategy to simulate more extreme settings. As shown in Fig. 5, we introduce a parameter $\Delta P$ to adjust image scales (equivalent to drone height), which can be expressed as: ($\Delta P$$>$0) means that making the height of drone-view seem higher (see Fig. 5(a)), while ($\Delta P$$<$0) means that making the height of drone-view seem lower (see Fig. 5(b)). $|\Delta P|$ denotes the degree of scale adjustment. Then, we roughly estimate the corresponding new height of synthesized images with $\Delta P$ as:
$H_{D}^{\prime}=H_{D}+\lambda_{aug}\cdot\Delta P,$ (15)
where $\lambda_{aug}$=0.7 is the empirical value, estimated by the visual scale and height of the existing drone-view images. By applying the aforementioned strategy to Long and Short sets, we generate test sets with extreme scales, named as Long ($\Delta P$$>$0) and Short ($\Delta P$$<$0), respectively. Note that drone-views within Long ($\Delta P$$>$0) exhibit higher heights (meaning smaller image scale), whereas those within Short ($\Delta P$$<$0) have lower heights (meaning larger image scale).
The SUES-200 [33] dataset collects drone-view images for 200 scenes at four heights (i.e., 150m, 200m, 250m, 300m) independently, where each scene contains 50 drone-view images and 1 satellite-view image. The training set includes 120 buildings, and the other 80 buildings are classified as the test set. The training and testing of models are performed with data of the same height. We highlight that despite the greater height differences of drone-captured images in SUES-200, SUES-200 exhibits weaker scale differences compared to University-1652 due to the different post-processing methods.
TABLE II: comparison with the state-of-the-art results reported on SUES-200. The input image size is 256$\times$256, the number of partitions $N$ is 4, and hyper-parameter $\alpha$ is 4.0. Best and second best performance are in red and blue colors.
Methods 150m 200m 250m 300m
R@1 AP R@1 AP R@1 AP R@1 AP
SUES-200 [33] 55.65 61.92 66.78 71.55 72.00 76.43 74.05 78.26
LCM [75] 43.42 49.65 49.42 55.91 54.47 60.31 60.43 65.78
LPN [25] 61.58 67.23 70.85 75.96 80.38 83.80 81.47 84.53
FSRA [20] 68.25 73.45 83.00 85.99 90.68 92.27 91.95 93.46
SaLPN 81.92 85.23 87.90 90.02 92.78 94.24 93.03 94.37
IV-B Implementation Details
@@
Model Settings. For the SaLPN framework, we set the number of partitions to $N$=4, and partition adjustment factor $\alpha$=14.0 by default. For SRGS, we adopt Chebyshev distance to generate coordinate map $CM$, and set the heatmap segmentation threshold $\delta$=0.5. For the University-1652 dataset, we roughly set drone-views height $H_{D}\in[123.5,256]$ and satellite-views height $H_{S}$=$189.75m$, respectively.
@@
Training Details. By default, we resize each input image to 512$\times$512 for both training and testing phases, and employ the ResNet-50 [73] to extract visual features. In training, we set the initial learning rate of 0.0001 for the backbone, and the rest of learnable parameters are set to 0.001. Stochastic gradient descent (SGD) is adopted with batch size 4, momentum 0.9, and weight decay 0.0005 for optimization. The model is trained for 120 epochs, and the learning rate is decayed by 0.1 after 80 epochs. We also employ random horizontal image flipping as data augmentation. During testing phase, Euclidean distance is adopted to measure the similarity between the query images and candidate images in the gallery set.
@@
Figure 6: Image retrieval results obtained with SaLPN, LPN and ResNet. Specifically, we present the correspondingly numbered image retrieval outcomes from the original test set (a), Long ($\Delta P$=100) (b), and Short ($\Delta P$=-100) (c), respectively.
IV-C Comparison with the State-of-the-art
@@
Performance on University-1652. As shown in Table I, we compare SaLPN with other methods on University-1652. Except for the standard test set provided, we employ Eq. (15) to generate two supplementary test sets, named Long ($\Delta P$=100) and Short ($\Delta P$=-100), which exhibit more extreme height of drone-views. In other words, Long ($\Delta P$=100) and Short ($\Delta P$=-100) have more significant cross-view scale differences. The quantitative results are divided into two groups, ResNet-based and ViT-based methods. On the ResNet-based track, SaLPN achieves 84.37% R@1 and 86.51% AP on default test set, 64.89% R@1 and 68.86% AP on Short ($\Delta P$=-100) set and 64.08% R@1, and 68.75% AP on Long ($\Delta P$=100) set, respectively. On the default test set, SaLPN achieves competitive performance with recent SDPL [53] and clearly outperforms other solutions, such as $D^{2}$-GeM [12] and Sample4Geo [37]. Moreover, it is remarkable that SaLPN yields superior results on both the Short ($\Delta P$=-100) and Long ($\Delta P$=100) compared with partition-based LPN [20] and SDPL [53]. Specifically, compared with LPN, SaLPN improves R@1 from 47.24%, 35.57% to 64.89% (+17.65%), 64.08% (+28.51%), respectively. On the ViT-based track, SaLPN is compatible with the ViT framework and has surpassed the sate-of-the-art methods on Short ($\Delta P$=-100) and Long ($\Delta P$=100), achieving an average R@1 and AP of 79.69% and 82.46%. Compared with Instance Loss [8] and FSRA, SaLPN significantly improves retrieval accuracy in all scenarios, with an average R@1 improvement of 9.83% and 5.61%. This shows that SaLPN is robust to scenes with inconsistent cross-view scales.
Performance on SUES-200. To further validate the effectiveness of SaLPN, we report the results on SUES-200 [33]. As shown in Table II, SaLPN achieves competitive results under different height. Compared with FSRA [20] at four heights, SaLPN boosts the R@1 from 68.25%, 83.00%, 90.68%, 91.95% to 81.92% (+13.67%), 87.90% (+4.90%), 92.78% (+2.10%), 93.03% (+1.08%), respectively. Different from drone-view images with scales variations in University-1652, the training set in SUES-200 has a uniform scale at a specific height. Besides, we observe that the cross-view scale difference in SUES-200 is weaker than University-1652. Therefore, we only leverage HAAS in the testing phase and experimentally determine $\alpha$=4.0.
@@
Qualitative Results. Moreover, we visualize some retrieved results in Fig. 6. To prove the anti-scaling robustness of SaLPN, we provide same numbered results generated by baseline [73] and LPN [25] on different test sets. For the original test set, both LPN and SaLPN are able to retrieve the correct satellite labels. As shown in Fig. 6 (b) and (c), when the drone-view images undergo more drastic scale variations, it can be observed that SaLPN still achieves correct results. This benefits from the fact that SaLPN explicitly aligns contents of partition pairs, thus mining robust contextual information. For example, as shown in Group 3 of Fig. 6 (b), while the target in the satellite-view is discernible, the target size in the drone-view is notably smaller compared to that in the satellite-view. This discrepancy leads to a misalignment of features extracted by LPN with fixed templates, resulting in inaccurate retrieval results. On the other hand, as shown in Group 2 of Fig. 6 (c), drone-view with lower height implies less environmental information and larger scale differences of cross-view images. Compared to LPN, SaLPN can dynamically adjust partition sizes to boost the alignment of local features, but also refine part-level features, yielding accurate retrieval results.
@@
TABLE III: ablation study on robustness of SaLPN to scale-inconsistent scenarios. we reported the R@1 accuracy for baseline, LPN, FSRA, SDPL and SaLPN on various test sets. The experimental results are divided into two groups: ResNet50-based at the top and ViT-based at the bottom. best and second best performance are in red and blue colors, respectively.
@@
Figure 7: Performance degradation of different methods on Short ($\Delta P\leq 0$) and Long ($\Delta P\geq 0$). Vertical axis: magnitude of decrease in R@1. Horizontal axis: degree of height adjustment $|\Delta P|$. (a) and (b) show the results of ResNet-based methods, while (c) and (d) show the results of ViT-based methods.
IV-D Ablation Studies and Further Discussion
To verify the effectiveness of SaLPN, we conduct extensive ablation studies, including:
  • To validate the robustness of SaLPN in scale-inconsistent scenarios and its adaptability to diverse backbone networks;
  • To prove the effectiveness of the proposed square partition strategy, height-aware adjustment strategy and saliency-guided refinement strategy;
  • To explore the effect of hyper-parameters on model performance, including number of partitions $N$, partition adjustment factor $\alpha$ and input image size.
Robustness of SaLPN against scale-inconsistent scenarios. As a hard-partition-based method, our method can alleviate the dependence on scale consistency of cross-view images by dynamically adjusting the size of partition pairs. To fully verify the robustness of SaLPN against scale-inconsistent scenarios, we first synthesize a series of test sets with lower and higher height of drone-views, named Short ($\Delta P$$<$0) and Long ($\Delta P$$>$0), where a larger $|\Delta P|$ represents a larger drone-view scale variation. Then we compare SaLPN with representative partition-based methods, such as LPN [25] and FSRA [20]. The results are shown in Table III, divided into ResNet50-based track and ViT-based track. The performance degradation of each method on different datasets is shown in Fig. 7. On the ResNet50-based track, SaLPN shows the best retrieval accuracy, with significantly less performance degradation than other partition-based methods. In particular, compared with LPN and SDPL, SaLPN improves average R@1 from 56.53%, 65.02% to 70.66% (+14.13%, +5.64%), respectively. Furthermore, the integration of HAAS into SaLPN led to a notable improvement in retrieval accuracy in extreme scenarios, with a 5.42% improvement in the average R@1. This highlights that HAAS facilitates alignment of fine-grained features from multiple views. On the ViT-based track, SaLPN also achieves optimal retrieval accuracy across datasets of all scales, with an average R@1 of 78.97%. Compared with FSRA and LPN, SaLPN improves average R@1 by 5.11% and 3.88%. This fully validates the robustness of SaLPN to cross-view scale variations, as well as the compatibility with various backbone networks. Moreover, Fig. 7 proves that HAAS can alleviate the performance degradation of SaLPN caused by cross-view scale inconsistency without additional learnable parameters. It is worth noting that the performance degradation of ViT-based methods is generally lower than that of ResNet-based methods. We conclude that the ViT framework is capable of capturing long-range feature dependencies, thereby contributing to its robustness against scale changes. As depicted in Fig. 7 (c), in contrast to ResNet-based methods, the performance degradation of ViT methods is comparable in Short ($\Delta P$$<$0), whereas the baseline exhibits the minimal degradation. This phenomenon can be attributed to the robust feature extraction capability of the ViT, coupled with the scarcity of effective features in Short ($\Delta P$$<$0) scenarios.
@@ IV-D Ablation Studies and Further Discussion
Effectiveness of primary components. The core components of SaLPN are SPS, HAAS and SGRS. We construct five derived models to validate the effectiveness of each strategy, and results are shown in Table IV. Firstly, fine-grained partitions generated by SPS contain both local and global features, thus learning rich contextual information. Compared to the baseline, SPS provides +19.76% R@1, +21.60% R@1 and +24.88% R@1 on three test sets. When HAAS is combined with SPS, the model’s accuracy is notably enhanced in scenarios involving scale inconsistency, while maintaining stable performance in the standard test set. For instance, after removing HAAS from SaLPN, the R@1 of model (SPS+SGRS) decreases by 0.83%, 4.81%, and 12.95% on standard test set, Short ($\Delta P$=-100) and Long ($\Delta P$=100), respectively. We emphasize that drone-views with low height contain incomplete depictions of targets and limited environmental information. consequently, it is plausible that the performance gain of HAAS is lower in the Short set compared to the Long set. To further improve the retrieval accuracy of SaLPN, we introduce SGRS to refine salient regions of partitions. It can be observed that SGRS boosts the retrieval accuracy of SaLPN in various scenarios, yielding 3.68%, 5.14% and 3.37% R@1 improvement over the model (SPS+HAAS).
@@
TABLE IV: performance comparison of SaLPN framework with different components. $\checkmark$ indicates that the module is applied.
Model structure Standard Test set Short ($\Delta P$=-100) Long ($\Delta P$=100)
SPS HAAS SGRS R@1 AP R@1 AP R@1 AP
@@
$\checkmark$ $\checkmark$ $\checkmark$ 84.37 86.51 64.89 68.86 64.08 68.75
Variants of SGRS. As mentioned above, SGRS segments part-level features based on heatmaps to obtain global features, salient features and background features, which are matched independently. To fully verify the effectiveness of SGRS, we construct a variety of SGRS variants, including different output combinations as well as coordinate map $CM$, as shown in Fig. 8. Table V shows the experimental results. Regardless of the specific combination of output modes employed by SGRS to refine partitions, it consistently outperforms the model without SGRS (i.e., only $g_{i}$). On the other hand, compared to Euclidean distance and Manhattan distance, the utilization of Chebyshev distance within SGRS yields an optimal gain, surpassing that of Manhattan distance by a notable margin. This phenomenon can be attributed to the fact that $CM$ generated by the Chebyshev paradigm exhibits a structural similarity to the square partition strategy employed. Therefore, it is within the realm of expectation that the model based on Euclidean distance yields comparable results. Since the feature segmentation process of SGRS is unsupervised, the injected $CM$ can assist SGRS in judging the importance of features and classifying neighborhood features into the same class. After removing of $CM$ from SGRS, the average R@1 and AP metrics of SaLPN decreased by 6.48% and 5.83%.
@@
TABLE V: ablation study on different variants of SGRS. The experiment aims to explore the impact of different feature combinations and coordinate maps on performance.
@@
Partition adjustment factor $\alpha$. Referring to Eq. (4), HAAS can adjust the degree of variation in partition size by modifying $\alpha$. An improperly choice for $\alpha$ may result in a misalignment of content between cross-view partition pairs. To ensure that $\alpha$ satisfies all test scenarios, $H_{D}^{max}$ and $H_{D}^{min}$ in Eq. 6 and 7 are from Long ($\Delta P$=+150) and Short ($\Delta P$=-150), respectively. Then, we set a series of $\alpha$ for experiments, and the results are shown in Table VI. It can be observed that when $\alpha$ is small, the size of partition pairs of SaLPN is almost of the same size, yielding favorable outcomes in the Middle set but suffering from significant performance degradation in extreme scenarios. When $\alpha$ increases from 0.0 to 14.0, the retrieval accuracy of SaLPN continues to rise in the Short ($\Delta P\leq 0$) and Long ($\Delta P\geq 0$) sets. In particular, the R@1 of our method improves from 17.84% and 34.93% to 25.68%(+7.84%) and 55.48%(+20.55%) on Short ($\Delta P$=-150) set and Long ($\Delta P$=+150) set, respectively. The average R@1 of SaLPN also increases from 63.83% to 70.66%(+6.83%). This verifies that HAAS can enhance the robustness of our model against scale changes. As $\alpha$ continues to increase, the retrieval accuracy of SaLPN exhibits a slight decline across various Short ($\Delta P\leq 0$) scenes. Considering the average performance of the model, we set $\alpha$=14.0 as the default setting.
TABLE VI: performance comparison of SaLPN with various partition adjustment factor $\alpha$. we reported the R@1 accuracy on three types of datasets, including Short ($\Delta P\leq 0$) set, Middle set and Long ($\Delta P\geq 0$) set.